FUNDAMENTALS AND TORSIONAL COMBINATION BANDS OF TWO ISOMERS OF THE OCS-CO_{2} COMPLEX

\underline{J. NOROOZ OLIAEE}, M. DEHGHANY, F. MIVEHVAR, MAHIN AFSHARI, N. MOAZZEN-AHMADI, Department of Physics and Astronomy, University of Calgary, 2500 University Dr., N.W., Calgary, Alberta T2N 1N4, Canada; A.R.W. McKELLAR, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, Ontario, K1A 0R6, Canada.

Infrared spectra of two isomers of the weakly bound OCS-CO_{2} complex were observed using a tunable diode laser to probe a pulsed supersonic slit-jet. Spectra were recorded in the regions of OCS ν_{1} (∼2060 cm^{−1}) and CO_{2} ν_{3} (∼2349 cm^{−1}) fundamental stretching vibrations. The lowest energy isomer (isomer a) was previously studied by microwave spectroscopy.^{a} Here we report the first infrared observation of isomer a and also detection of a new higher energy form (isomer b). Structures were determined with the help of isotopic substitution. Both isomers are planar, with slipped near-parallel geometries. In isomer a, the intermolecular (centre of mass) separation is 3.55 Å and the C atom of the CO_{2} is closer to the S atom of the OCS. In isomer b, the C atom of CO_{2} slides closer to the O atom of OCS and the center of mass separation increases to 3.99 Å.^{b} Three combination bands involving the intermolecular torsional (out-of-plane bend) vibrations were also analyzed. The out-of-plane torsional frequencies were measured to be 18.8 cm^{−1} and 15.9 cm^{−1} for isomers a and b, respectively, indicating that isomer a is indeed more bound than isomer b.

aStewart E. Novick, R. D. Seunram and F. J. Lovas, J. Chem. Phys. 88, 687-690 (1988).

bM. Dehghany, J. Norooz Oliaee, Mahin Afshari, N. Moazzen-Ahmadi and A. R. W. McKellar, J. Chem. Phys. 130, 224310